# CALF: Aligning LLMs for Time Series Forecasting via Cross-modal Fine-Tuning

### v2

@@ Abstract
Deep learning (e.g., Transformer) has been widely and successfully used in multivariate time series forecasting (MTSF). Unlike existing methods that focus on training models from a single modal of time series input, large language models (LLMs) based MTSF methods with cross-modal text and time series input have recently shown great superiority, especially with limited temporal data. However, current LLM-based MTSF methods usually focus on adapting and fine-tuning LLMs, while neglecting the distribution discrepancy between textual and temporal input tokens, thus leading to sub-optimal performance. To address this issue, we propose a novel Cross-ModAl LLM Fine-Tuning (CALF) framework for MTSF by reducing the distribution discrepancy between textual and temporal data, which mainly consists of the temporal target branch with temporal input and the textual source branch with aligned textual input. To reduce the distribution discrepancy, we develop the cross-modal match module to first align cross-modal input distributions. Additionally, to minimize the modality distribution gap in both feature and output spaces, feature regularization loss is developed to align the intermediate features between the two branches for better weight updates, while output consistency loss is introduced to allow the output representations of both branches to correspond effectively. Thanks to the modality alignment, CALF establishes state-of-the-art performance for both long-term and short-term forecasting tasks with low computational complexity, and exhibiting favorable few-shot and zero-shot abilities similar to that in LLMs. Code is available at https://github.com/Hank0626/CALF.
1 Introduction
@@
Multivariate time series forecasting (MTSF) plays a crucial role in the domain of time series analysis and has further boasted a wide range of real-world applications including weather forecasting [1], energy prediction [2], financial modeling [3]. To achieve more accurate forecasting performance, numerous deep learning-based MTSF methods trained on a single modal of time series input have been developed in recent years [4, 5, 6, 7, 8, 9, 10, 11] and have gained great success.
@@
However, previous single-modal MTSF methods [12] may suffer from overfitting problems, due to the limited training data, thus limiting their real applications. To relieve such issues, some pioneering works attempt to introduce the powerful Large Language Models (LLMs) models in time series forecasting by employing the strong context modeling ability of LLMs. For example, Zhou et al. [13] proposed a unified time series analysis framework by adapting and fine-tuning LLMs. Building upon this, other works have introduced additional enhancements to further expand the capabilities of LLMs in time series forecasting, including refining fine-tuning methods [14], sequence decomposition [15], and the incorporation of textual prompts [12]. Benefiting from the large-scale pre-training, LLM-based methods not only exhibit strong context modeling capabilities but also help mitigate the problem of overfitting.
Despite the great success of LLM-based MTSF methods, existing LLM-based MTSF methods usually focus on adapting and fine-tuning LLMs, while neglecting the distribution discrepancy between textual and temporal input tokens, thus leading to sub-optimal performance. In practice, current LLM-based methods typically treat pre-trained LLMs as well-initialized forecasting models and project time series data using a simple linear layer as input for the LLMs. While this straightforward approach is intuitive, it can lead to sub-optimal results due to significant distribution discrepancies between textual and temporal data. As shown in Fig. 1(a), we show the distribution of textual and temporal tokens of LLM-based MTSF methods, and we find that the temporal tokens in existing LLM-based methods cannot align well with the original textual tokens from LLMs [13, 16, 12, 14]. These observations inspire us to develop a Cross-modal LLM Fine-Tuning framework to consider the distribution discrepancy between textual and temporal input tokens.
Figure 1: (a) The t-SNE visualization of pre-trained word token embeddings of LLM with temporal tokens of ETTh2 dataset from GPT4TS [13] (Left) and our method (Right). Our method shows more cohesive integration, indicating effective modality alignment. Appendix A shows more results. (b) Conceptual illustration of cross-modal fine-tuning technique.
@@
Inspired by the above observations, we propose a Cross-ModAl LLM Fine-Tuning (CALF) framework, which employs cross-modal fine-tuning to allow more comprehensive alignment between temporal target modalities and textual source modalities. Specifically, CALF consists of two branches: the temporal target branch and the textual source branch. The temporal target branch processes time series information, while the textual source branch extracts and adapts information from pre-trained LLMs using aligned textual modal tokens. To bridge the modality gap between these branches, we introduce three meticulously designed cross-modal fine-tuning techniques (see Fig. 1(b)): (1) Cross-Modal Match Module integrates time series and textual inputs through principal word embedding extraction and a cross-attention mechanism, ensuring efficient alignment of the marginal input distribution between time series and text; (2) Feature Regularization Loss aligns the outputs of each intermediate layer, ensuring that gradients at every layer are more effectively guided for better weight updates; (3) Output Consistency Loss ensures that the output representations of textual and temporal series modalities correspond effectively, resolving discrepancies in the representation space and maintaining consistent semantic context for time series data. Through a more comprehensive alignment, our CALF consistently achieves state-of-the-art performance in both long-term and short-term forecasting across multiple datasets, demonstrating excellent few/zero-shot generalization capabilities, while maintaining significantly low complexity.
@@
The contributions of this paper are threefold: (i) We identify the significant distribution discrepancies between textual and temporal modalities in existing LLM-based forecasting models and highlight the importance of addressing this misalignment for improved performance. (ii) We propose CALF, a novel framework that employs cross-modal fine-tuning techniques to comprehensively align temporal and textual data. The framework includes three specific methods: the Cross-Modal Match Module for aligning input distributions, Feature Regularization Loss for better gradient guidance and weight updates, and Output Consistency Loss for resolving output representation space discrepancies and maintaining consistent semantic context. (iii) Extensive experiments on eight real-world datasets demonstrate that CALF achieves state-of-the-art performance on both long-term and short-term time series forecasting tasks, with favorable generalization ability and low computational complexity.
2 Related Work
2.1 Time Series Forecasting
@@
In recent years, deep learning has significantly revolutionized the field of time series forecasting, with a plethora of methods emerging to enhance predictive accuracy [7, 8, 4, 17, 18, 19, 20]. Among these, Transformer-based models have emerged as the frontrunners, offering unparalleled performance due to their exceptional ability to model complex dependencies in data  [6, 5, 21, 22, 9, 11, 20]. However, they often have limitations due to the scarcity of training data, overfitting in specific domains, and the necessity for intricate architectural designs.
@@
In response to these challenges, the integration of LLMs into time series forecasting has emerged as a novel and promising direction. This approach leverages the extensive pre-training of LLMs to enhance the context-modeling capacity in time series analysis. A groundbreaking framework proposed by Zhou et al. [13] first demonstrated the potential of adapting LLMs for time series analysis. Following this paradigm, subsequent research has introduced further refinements and innovations. For example, Chang et al. [14] introduced a novel two-stage fine-tuning method and integrated time-series patching with additional temporal encoding into pre-trained LLMs. Cao et al. [15] incorporated decomposition of time series and selection-based prompts for adapting to non-stationary data. However, these works often directly input time series data into LLMs, overlooking the misalignment between time series and textual modalities. Some works have attempted to address this issue. Sun et al. [16] aligned time series data with LLM embeddings using contrastive learning and employed soft prompts for effective time series task handling. Jin et al. [12] reprogrammed time series input with text prototypes and enriches it using context as a prefix for LLM alignment. Despite these efforts, the alignment strategies have not been sufficiently effective.
2.2 Cross-Modal Fine-tuning
@@
The objective of cross-modal fine-tuning is to apply models pre-trained on data-rich modalities to data-scarce modalities, addressing issues of data insufficiency and poor generalization [23]. Many existing works focus on transferring LLMs to other modalities, such as vision [24, 25], audio [26, 27], and biology [28, 29]. These efforts provide initial evidence of the cross-modal transfer capacity of pre-trained models. In the domain of time series, current research primarily leverages the powerful contextual modeling capabilities of LLMs to fine-tune them for improved forecasting performance [13, 12, 15, 14, 16], often neglecting the gap between the input and output distributions of language and time series modalities. In this work, we apply cross-modal fine-tuning techniques to address the challenge of transferring pre-trained language model knowledge to the time series modality.
3 Methodology
@@
3.1 Cross-Modal Match Module
As demonstrated in previous work [30], the matrices of word embedding layers in pre-trained LLMs constitute a well-structured context representation space, e.g., semantic distances between different words can be quantified through vector similarity. This word embedding layer represents the input distribution of the language modality in pre-trained LLMs. Despite this promising property, previous LLM-based time series methods often overlook this distribution, instead projecting the time series data to match the input dimensions of the language model [13, 15, 14].
In this work, we attempt to align the input distribution of time series with the word embedding of LLMs. Therefore, we propose a cross-modal match module to deal with this problem. Specifically, given a multivariate time series $I\in\mathbb{R}^{T\times C}$ as input, where $T$ is the input sequence length and $C$ is the number of variants, we first use the embedding layer similar to [31], followed by Multi-head Self Attention (MHSA) to get the projected time tokens $X_{time}$:
@@ 3.1 Cross-Modal Match Module
After that, we consider using cross-attention to align $X_{time}$ from the temporal modality and the word embedding dictionaries $\mathcal{D}\in\mathbb{R}^{|\mathcal{A}|\times M}$, where $|\mathcal{A}|$ is the size of the alphabet, to the textual modality. However, due to $|\mathcal{A}|$ is usually huge, e.g., 50257 in GPT2 [32], directly using cross-attention incurs significant cost. Observing that semantic-similar words form “synonym clusters”, we propose a principal word embedding extraction strategy, which uses the cluster center to represent surrounding words, to reduce the number of word entries. Specifically, we use Principal Component Analysis (PCA) to perform dimension reduction on $\mathcal{D}$ to obtain the principal word embeddings $\hat{\mathcal{D}}\in\mathbb{R}^{d\times M}$,
@@
3.2 Feature Regularization Loss
@@
$\mathcal{L}_{feature}=\sum_{i=1}^{L}\gamma^{(L-i)}{\rm sim}(\phi^{text}_{i}(F_{{text}}^{l}),\phi^{time}_{i}(F_{{time}}^{l})),$ (4)
where $\gamma$ is a hyper-parameter that controls the loss scale from different layers, and ${\rm sim}(\cdot,\cdot)$ is a chosen similarity function, such as $L_{1}$ loss. Following [33], we introduce two trainable projection layers $\phi^{text}_{l}(\cdot)$ and $\phi^{time}_{l}(\cdot)$ to transform the features from textual and temporal modalities to the shared representation space.
3.3 Output Consistency Loss
@@
3.4 Parameter Efficient Training
To avoid catastrophic forgetting and improve training efficiency, we employ the parameter-efficient training technique to fine-tune the pre-trained LLMs. Specifically, for the temporal target branch, we introduce Low-rank Adaptation (LoRA) [34] and fine-tune the positional encoding weights. The total loss during training is the weighted summation of the supervised loss $\mathcal{L}_{sup}$, the feature regularization loss $\mathcal{L}_{feature}$, and the output consistency loss $\mathcal{L}_{output}$:
@@ 3.4 Parameter Efficient Training
4 Experiments
To demonstrate the effectiveness of the proposed CALF, we conduct extensive experiments on various time series forecasting tasks, including long/short-term forecasting and few/zero-shot learning. Additionally, we validate the model with low complexity, highlighting its efficiency in practical applications.
Baselines.
We carefully select representative baselines from the recent time series forecasting landscape, including the following categories: (1) LLMs-based models: TimeLLM [12] and GPT4TS [13]; (2) Transformer-based models: PatchTST [6], iTransformer [31], Crossformer [5], ETSformer [21], FEDformer [9] and Autoformer [22]; (3) CNN-based models: TCN [35], MICN [17] and TimesNet [4]; (4) MLP-based models: DLinear [7] and TiDE [8]. Besides, N-HiTS [36] and N-BEATS [37] are included for short-term forecasting.
Implementation Details.
Following [13], we use pre-trained GPT2 based model [32] with the first 6 Transformer layers as our backbone. Optimization is conducted using the Adam optimizer [38], with a learning rate of $0.0005$. For the total loss function, we set the hyper-parameters $\gamma=0.8$, $\lambda_{1}=1$ and $\lambda_{2}=0.01$. In terms of loss functions for long-term forecasting, we apply L1 loss across all three loss types for ETT datasets, while for the other three datasets, smooth L1 loss is utilized. For short-term forecasting, we compute supervised loss with SMAPE, modal consistency loss with MASE, and feature regularization loss with smooth L1 loss, respectively. More details are provided in Appendix D.
4.1 Long-term Forecasting
Setups.
We conduct experiments on seven widely-used real-world datasets, including the Electricity Transformer Temperature (ETT) dataset with its four subsets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, and Traffic [22]. Detailed descriptions of datasets are provided in Sec. C.1. The input time series length $T$ is fixed as $96$ for a fair comparison, and we adopt four distinct prediction horizons $H\in\{96,192,336,720\}$. Consistent with prior works, the Mean Square Error (MSE) and Mean Absolute Error (MAE) are chosen as evaluation metrics.
Results.
Comprehensive long-term forecasting results are presented in Tab. 1. Our method consistently delivers state-of-the-art performance, achieving the top results in 56 evaluations, in contrast to the nearest competing baseline which achieves top results only 7 times. Notably, our approach reduces MSE/MAE by 7.05%/6.53% compared to the state-of-the-art Transformer-based model PatchTST. In comparison with the LLM-powered method GPT4TS, we observe a reduction of 5.94%/5.14% in MSE/MAE. Moreover, our improvements are substantial against other baseline methods, exceeding 10% in most cases.
Models CALF TimeLLM GPT4TS PatchTST iTransformer Crossformer FEDformer TimesNet MICN DLinear TiDE
[12]
[13]
[6]
[31]
[5]
[9]
[4]
[17]
[7]
[8]
@@ Results.
    We utilize their official codebase with the same experimental setup as ours, including input length and a GPT2 model with 6 layers, to ensure the fairness of the results. Other results are obtained from [31].
Table 1: Multivariate long-term forecasting results. The input sequence length $T$ is set to 96 for all baselines. All the results are averaged from 4 different prediction lengths $H\in\{96,192,336,720\}$. The best and second best results are in bold and underlined. Sec. F.1 shows the full results.
4.2 Short-term Forecasting
Setups.
We adopt the M4 datasets [39], which comprise univariate marketing data collected yearly, quarterly, and monthly. Comprehensive details are available in Sec. C.2. In this case, the prediction horizons are comparatively short, ranging in $[6,48]$. Correspondingly, the input lengths are set to be twice the size of the prediction horizons. The evaluation metrics are symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MSAE), and overall weighted average (OWA).
Results.
As shown in Tab. 2, our method demonstrates superior performance in short-term forecasting across various evaluation metrics. Notably, it achieves the best results in 14 out of 15 categories, markedly outperforming all baselines. In comparison with TimesNet, currently the leading method in short-term forecasting, our model achieves a 1% overall improvement in performance.
Models CALF TimeLLM GPT4TS PatchTST ETSformer FEDformer Autoformer TimesNet TCN N-HiTS N-BEATS DLinear
[12]
[13]
[6]
[21]
[9]
[22]
[4]
[35]
[36]
[37]
[7]
@@
Average
SMAPE
11.765 11.983 11.991 12.059 14.718 12.840 12.909 11.829 13.961 11.927 11.851 13.639
MASE
1.567 1.595 1.600 1.623 2.408 1.701 1.771 1.585 1.945 1.613 1.599 2.095
OWA
0.844 0.859 0.861 0.869 1.172 0.918 0.939 0.851 1.023 0.861 0.855 1.051
Table 2: Short-term forecasting results on M4 dataset. The input length and prediction length are set to $[12,96]$ and $[6,48]$, respectively. Sec. F.2 shows the full results.
4.3 Few/zero-shot Learning
LLMs have demonstrated remarkable performance in both few-shot and zero-shot tasks. The capabilities of few-shot and zero-shot learning are critically important for general time series forecasting models [40, 41, 42, 43]. To thoroughly assess the generalized ability of our method in time series forecasting, we conduct experiments under few-shot and zero-shot learning settings. In few-shot learning, only a small ratio of the training data is utilized. For zero-shot learning, the model trained on one dataset is directly employed for testing on another dataset without any additional training.

### v3

@@ Abstract
Deep learning (e.g., Transformer) has been widely and successfully used in multivariate time series forecasting (MTSF). Unlike existing methods that focus on training models from a single modal of time series input, large language models (LLMs) based MTSF methods with cross-modal text and time series input have recently shown great superiority, especially with limited temporal data. However, current LLM-based MTSF methods usually focus on adapting and fine-tuning LLMs, while neglecting the distribution discrepancy between textual and temporal input tokens, thus leading to sub-optimal performance. To address this issue, we propose a novel Cross-ModAl LLM Fine-Tuning (CALF) framework for MTSF by reducing the distribution discrepancy between textual and temporal data, which mainly consists of the temporal target branch with temporal input and the textual source branch with aligned textual input. To reduce the distribution discrepancy, we develop the cross-modal match module to first align cross-modal input distributions. Additionally, to minimize the modality distribution gap in both feature and output spaces, feature regularization loss is developed to align the intermediate features between the two branches for better weight updates, while output consistency loss is introduced to allow the output representations of both branches to correspond effectively. Thanks to the modality alignment, CALF establishes state-of-the-art performance for both long-term and short-term forecasting tasks with low computational complexity, and exhibits favorable few-shot and zero-shot abilities similar to that in LLMs.
Codehttps://github.com/Hank0626/CALF
Introduction
@@
Multivariate time series forecasting (MTSF) plays a crucial role in the domain of time series analysis and has further boasted a wide range of real-world applications including weather forecasting (Angryk et al. 2020), energy prediction (Demirel et al. 2012), financial modeling (Patton 2013). To achieve more accurate forecasting performance, numerous deep learning-based MTSF methods trained on a single modal of time series input have been developed in recent years (Wu et al. 2023; Zhang and Yan 2023; Nie et al. 2023; Zeng et al. 2023; Zhou et al. 2022; Wen et al. 2022; Dai et al. 2024; Qiu et al. 2024b) and have gained great success.
@@
However, previous single-modal MTSF methods may suffer from overfitting problems, due to the limited training data, thus limiting their real applications. To relieve such issues, some pioneering works attempt to introduce the powerful Large Language Models (LLMs) models in time series forecasting by employing the strong context modeling ability of LLMs. For example, Zhou et al. (Zhou et al. 2023) proposed a unified time series analysis framework by adapting and fine-tuning LLMs. Building upon this, other works have introduced additional enhancements to further expand the capabilities of LLMs in time series forecasting, including refining fine-tuning methods (Chang, Peng, and Chen 2023), sequence decomposition (Cao et al. 2024), and the incorporation of textual prompts (Jin et al. 2024). Benefiting from the large-scale pre-training, LLM-based methods not only exhibit strong context modeling capabilities but also help mitigate the problem of overfitting.
Despite the great success of LLM-based MTSF methods, existing methods suffer unfavorable results on common MTSF benchmarks, as well as pool performance on few-shot or zero-shot tasks. As shown in Fig. 1, we visualize the distribution of the textual and temporal tokens of existing LLM-based MTSF methods, and it can be seen that the temporal tokens in existing LLM-based methods cannot align well with the original textual tokens from LLMs (Zhou et al. 2023; Sun et al. 2024; Jin et al. 2024; Chang, Peng, and Chen 2023). As a result, this distribution discrepancy contributes in part to the sub-optimal performance of the current approaches. We argue that the main reason for the above misalignment is due to the fact that existing methods only investigate the input side to match the feature dimensions between time series and LLM, and this simple solution hinders the manipulation of LLM’s powerful generalization capabilities. The above observations motivate us to develop a multi-level cross-modal aligning framework to alleviate the distribution discrepancy between temporal modal input and textural modal weights.
Figure 1: The t-SNE visualization of pre-trained word token embeddings of LLM with the hidden features from the penultimate layer from GPT4TS (Zhou et al. 2023), TimeLLM (Jin et al. 2024), TEST (Sun et al. 2024), and ours of ETTh2 dataset. Current LLM-based methods either use linear layers to project time series to the LLM’s feature dimension (Zhou et al. 2023) or employ cross-attention and contrastive learning techniques (Jin et al. 2024; Sun et al. 2024), which address only the input side and overlook alignment in the deeper layers. Our CALF achieves better alignment through multi-level cross-modal fine-tuning.
@@
Inspired by the above observations, we propose a Cross-ModAl LLM Fine-Tuning (CALF) framework, which employs cross-modal fine-tuning to allow more comprehensive alignment between temporal target modalities and textual source modalities. Specifically, CALF consists of two branches: the temporal target branch and the textual source branch. The temporal target branch processes time series information, while the textual source branch extracts and adapts information from pre-trained LLMs using aligned textual modal tokens. To bridge the modality gap between these branches, we introduce three meticulously designed cross-modal fine-tuning techniques (see Fig. 2): (1) Cross-Modal Match Module integrates time series and textual inputs through principal word embedding extraction and a cross-attention mechanism, ensuring efficient alignment of the marginal input distribution between time series and text; (2) Feature Regularization Loss aligns the outputs of each intermediate layer, ensuring that gradients at every layer are more effectively guided for better weight updates; (3) Output Consistency Loss ensures that the output representations of textual and temporal series modalities correspond effectively, resolving discrepancies in the representation space and maintaining consistent semantic context for time series data. Through a more comprehensive alignment, our CALF consistently achieves state-of-the-art performance in both long-term and short-term forecasting across multiple datasets, demonstrating excellent few/zero-shot generalization capabilities, while maintaining significantly low complexity.
The contributions of this paper are threefold:
  1. 1.
    We identify the significant distribution discrepancies between textual and temporal modalities in existing LLM-based forecasting models and highlight the importance of addressing this misalignment for improved performance.
  2. 2.
@@
    We propose CALF, a novel framework that employs cross-modal fine-tuning techniques to comprehensively align temporal and textual data. The framework includes three specific methods: the Cross-Modal Match Module for aligning input distributions, Feature Regularization Loss for better gradient guidance and weight updates, and Output Consistency Loss for resolving output representation space discrepancies and maintaining consistent semantic context.
  3. 3.
    Extensive experiments on eight real-world datasets demonstrate that CALF achieves state-of-the-art performance on both long-term and short-term time series forecasting tasks, with favorable generalization ability and low computational complexity.
Related Work
Time Series Forecasting
@@
In recent years, deep learning has significantly revolutionized the field of time series forecasting, with a plethora of methods emerging to enhance predictive accuracy (Zeng et al. 2023; Das et al. 2023; Wu et al. 2023; Wang et al. 2022; Liu et al. 2023a). Among these, Transformer-based models have emerged as the frontrunners, offering unparalleled performance due to their exceptional ability to model complex dependencies in data  (Nie et al. 2023; Zhang and Yan 2023; Woo et al. 2022; Wu et al. 2021; Zhou et al. 2022; Dai et al. 2024). However, they often have limitations due to the scarcity of training data, overfitting in specific domains, and the necessity for intricate architectural designs.
@@
To tackle these challenges, the integration of LLMs into time series forecasting has emerged as a novel and promising direction. This approach leverages the extensive pre-training of LLMs to enhance the context-modeling capacity in time series analysis. A groundbreaking framework proposed by Zhou et al. (Zhou et al. 2023) first demonstrated the potential of adapting LLMs for time series analysis. Following this paradigm, subsequent research has introduced further refinements and innovations. For example, Chang et al. (Chang, Peng, and Chen 2023) introduced a novel two-stage fine-tuning method and integrated time-series patching with additional temporal encoding into pre-trained LLMs. Cao et al. (Cao et al. 2024) incorporated decomposition of time series and selection-based prompts for adapting to non-stationary data. However, these works often directly input time series data into LLMs, overlooking the misalignment between time series and textual modalities. Some works have attempted to address this issue. Sun et al. (Sun et al. 2024) aligned time series data with LLM embeddings using contrastive learning and employed soft prompts for effective time series task handling. Jin et al. (Jin et al. 2024) reprogrammed time series input with text prototypes and enriches it using context as a prefix for LLM alignment. Despite these efforts, the alignment strategies have not been sufficiently effective.
Cross-Modal Fine-tuning
@@
The objective of cross-modal fine-tuning is to apply models pre-trained on data-rich modalities to data-scarce modalities, addressing issues of data insufficiency and poor generalization (Shen et al. 2023). Many existing works focus on transferring LLMs to other modalities, such as vision (Kiela et al. 2019; Verma et al. 2024), audio (Jin et al. 2023; Hassid et al. 2024), and biology (Vinod, Chen, and Das 2023; Xiao et al. 2021). These efforts provide initial evidence of the cross-modal transfer capacity of pre-trained models. In the domain of time series, current research primarily leverages the powerful contextual modeling capabilities of LLMs to fine-tune them for improved forecasting performance (Zhou et al. 2023; Jin et al. 2024; Cao et al. 2024; Chang, Peng, and Chen 2023; Sun et al. 2024), often neglecting the gap between the input and output distributions of language and time series modalities. In this work, we apply cross-modal fine-tuning techniques to address the challenge of transferring pre-trained language model knowledge to the time series modality.
Methodology
@@
Cross-Modal Match Module
As demonstrated in previous work (Mikolov et al. 2013), the matrices of word embedding layers in pre-trained LLMs constitute a well-structured context representation space, e.g., semantic distances between different words can be quantified through vector similarity. This word embedding layer represents the input distribution of the language modality in pre-trained LLMs. Despite this promising property, previous LLM-based time series methods often overlook this distribution, instead projecting the time series data to match the input dimensions of the language model (Zhou et al. 2023; Cao et al. 2024; Chang, Peng, and Chen 2023).
In this work, we attempt to align the input distribution of time series with the word embedding of LLMs. Therefore, we propose a cross-modal match module to deal with this problem. Specifically, given a multivariate time series $I\in\mathbb{R}^{T\times C}$ as input, where $T$ is the input sequence length and $C$ is the number of variants, we first use the embedding layer similar to (Liu et al. 2024), followed by Multi-head Self Attention (MHSA) to get the projected time tokens $X_{time}$:
@@
After that, we consider using cross-attention to align $X_{time}$ from the temporal modality and the word embedding dictionaries $\mathcal{D}\in\mathbb{R}^{|\mathcal{A}|\times M}$, where $|\mathcal{A}|$ is the size of the alphabet, to the textual modality. However, due to $|\mathcal{A}|$ is usually huge, e.g., 50257 in GPT2 (Radford et al. 2019), directly using cross-attention incurs significant cost. Observing that semantic-similar words form “synonym clusters”, we propose a principal word embedding extraction strategy, which uses the cluster center to represent surrounding words, to reduce the number of word entries. Specifically, we use Principal Component Analysis (PCA) to perform dimension reduction on $\mathcal{D}$ to obtain the principal word embeddings $\hat{\mathcal{D}}\in\mathbb{R}^{d\times M}$,
@@
Feature Regularization Loss
@@
$\mathcal{L}_{feature}=\sum_{l=1}^{L}\gamma^{(L-l)}{\rm sim}(\phi^{text}_{l}(F_{{text}}^{l}),\phi^{time}_{l}(F_{{time}}^{l})),$ (4)
where $\gamma$ is a hyper-parameter that controls the loss scale from different layers, and ${\rm sim}(\cdot,\cdot)$ is a chosen similarity function, such as $L_{1}$ loss. Following (Chen et al. 2020), we introduce two trainable projection layers $\phi^{text}_{l}(\cdot)$ and $\phi^{time}_{l}(\cdot)$ to transform the features from textual and temporal modalities to the shared representation space.
Output Consistency Loss
@@
Parameter Efficient Training
To avoid catastrophic forgetting and improve training efficiency, we employ the parameter-efficient training technique to fine-tune the pre-trained LLMs. Specifically, for the temporal target branch, we introduce Low-rank Adaptation (LoRA) (Hu et al. 2021) and fine-tune the positional encoding weights. The total loss during training is the weighted summation of the supervised loss $\mathcal{L}_{sup}$, the feature regularization loss $\mathcal{L}_{feature}$, and the output consistency loss $\mathcal{L}_{output}$:
@@
Experiments
To demonstrate the effectiveness of the CALF, we conduct extensive experiments on various time series forecasting tasks, including long/short-term forecasting and few/zero-shot learning. Besides, we validate the model with low complexity, highlighting its efficiency in practical applications.
Models
CALF TimeLLM GPT4TS PatchTST iTransformer Crossformer FEDformer TimesNet MICN DLinear TiDE
@@
    We utilize their official codebase with the same experimental setup as ours, including input length and a GPT2 model with 6 layers, to ensure the fairness of the results. Other results are obtained from (Liu et al. 2024).
Table 1: Multivariate long-term forecasting results. The input sequence length $T$ is set to 96 for all baselines. The best and second best results are in bold and underlined.
Models
CALF TimeLLM GPT4TS PatchTST ETSformer FEDformer Autoformer TimesNet TCN N-HiTS N-BEATS DLinear
@@
Table 2: Short-term forecasting results on M4 dataset. The input length and prediction length are set to $[12,96]$ and $[6,48]$, respectively.
Baselines.
We carefully select representative baselines from the recent time series forecasting landscape, including the following categories: (1) LLMs-based models: TimeLLM (Jin et al. 2024) and GPT4TS (Zhou et al. 2023); (2) Transformer-based models: PatchTST (Nie et al. 2023), iTransformer (Liu et al. 2024), Crossformer (Zhang and Yan 2023), ETSformer (Woo et al. 2022), FEDformer (Zhou et al. 2022) and Autoformer (Wu et al. 2021); (3) CNN-based models: TCN (Bai, Kolter, and Koltun 2018), MICN (Wang et al. 2022) and TimesNet (Wu et al. 2023); (4) MLP-based models: DLinear (Zeng et al. 2023) and TiDE (Das et al. 2023). Besides, N-HiTS (Challu et al. 2022) and N-BEATS (Oreshkin et al. 2019) are included for short-term forecasting.
Implementation Details.
Following (Zhou et al. 2023), we use pre-trained GPT2 based model (Radford et al. 2019) with the first 6 Transformer layers as our backbone. Optimization is conducted using the Adam optimizer (Kingma and Ba 2014), with a learning rate of $0.0005$. For the total loss function, we set the hyper-parameters $\gamma=0.8$, $\lambda_{1}=1$ and $\lambda_{2}=0.01$. In terms of loss functions for long-term forecasting, we apply L1 loss across all three loss types for ETT datasets, while for the other three datasets, smooth L1 loss is utilized. For short-term forecasting, we compute supervised loss with SMAPE, modal consistency loss with MASE, and feature regularization loss with smooth L1 loss, respectively.
Long-term Forecasting
Setups.
We conduct experiments on seven widely-used real-world datasets, including the Electricity Transformer Temperature (ETT) dataset with its four subsets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, ECL, and Traffic (Wu et al. 2021). The input time series length $T$ is fixed as $96$ for a fair comparison, and we adopt four distinct prediction horizons $H\in\{96,192,336,720\}$. Consistent with prior works (Li et al. 2024), the Mean Square Error (MSE) and Mean Absolute Error (MAE) are chosen as evaluation metrics. Following TBF’s (Qiu et al. 2024a) setting, we do not use the drop last trick to ensure fair comparison.
Results.
Comprehensive long-term forecasting results are presented in Tab. 1, where all the results are averaged from different prediction lengths. Notably, our approach reduces MSE/MAE by 7.05%/6.53% compared to the state-of-the-art Transformer-based model PatchTST. In comparison with the LLM-powered method TimeLLM, we observe a reduction of 5.98%/5.34% in MSE/MAE. Moreover, our improvements are substantial against other baseline methods, exceeding 10% in most cases.
Short-term Forecasting
Setups.
We adopt the M4 datasets (Makridakis, Spiliotis, and Assimakopoulos 2018), which comprise univariate marketing data collected yearly, quarterly, and monthly. In this case, the prediction horizons are comparatively short, ranging in $[6,48]$. Correspondingly, the input lengths are set to be twice the size of the prediction horizons. The evaluation metrics are symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MSAE), and overall weighted average (OWA).
Results.
As shown in Tab. 2, our method demonstrates superior performance in short-term forecasting across various evaluation metrics. In comparison with TimesNet, currently the leading method in short-term forecasting, our model achieves a 1% overall improvement in performance.
Few/zero-shot Learning
LLMs have demonstrated remarkable performance in both few-shot and zero-shot tasks. The capabilities of few-shot and zero-shot learning are critically important for general time series forecasting models (Brown et al. 2020; Liu et al. 2023b; Kojima et al. 2022). To thoroughly assess the generalized ability of our method in time series forecasting, we conduct experiments under few-shot and zero-shot learning settings. In few-shot learning, only a small ratio of the training data is utilized. For zero-shot learning, the model trained on one dataset is directly employed for testing on another dataset without any additional training. All the results are averaged from 4 different prediction lengths $H\in\{96,192,336,720\}$.
Models ETTm1 ETTm2 ETTh1 ETTh2
MSE MAE MSE MAE MSE MAE MSE MAE
TiDE 0.515 0.469 0.303 0.337 0.779 0.604 0.421 0.428
DLinear 0.567 0.499 0.329 0.382 0.647 0.552 0.441 0.458
MICN 0.970 0.674 1.073 0.716 1.405 0.814 2.533 1.158
TimesNet 0.673 0.534 0.321 0.354 0.865 0.625 0.476 0.463
FEDformer 0.696 0.572 0.356 0.392 0.750 0.607 0.553 0.525
Crossformer 1.340 0.848 1.985 1.048 1.744 0.914 3.139 1.378
PatchTST 0.557 0.483 0.295 0.334 0.683 0.546 0.550 0.487
GPT4TS 0.608 0.500 0.303 0.336 0.689 0.555 0.579 0.497
TimeLLM 0.636 0.512 0.348 0.343 0.765 0.584 0.589 0.498
CALF 0.504 0.462 0.302 0.330 0.644 0.541 0.419 0.427
Table 3: Few-shot learning results on 10% training data of ETT datasets.